# Training-free LLM Verification via Recycling Few-shot Examples

### v1

@@ B.3 Additional comparison with few-shot prompting-based selection methods
As shown in Table 8, ReFeri consistently achieves the best or second-best accuracy across all LLMs and benchmarks. Interestingly, we observe that adding few-shot demonstrations to USC often degrades performance (e.g., on GPQA and DROP with GPT-4o-mini and LLaMA-3.1-8B), which is likely due to the sensitivity of LLMs to prompt format and positional bias of the responses. Notably, we observe that both prompt-based selection methods, USC and LLM-as-Judge, are highly sensitive to the order of candidate responses. In our experiments, over 90% of USC selections were made from the first two responses regardless of correctness (see Table 9). This highlights a critical weakness in prompt-based selection: the output is often determined more by position than content. In contrast, our approach mitigates such ordering artifacts by decoupling few-shot demonstrations from the selection prompt and using them only for scoring. Furthermore, LLM-as-Judge does not perform reliably on more complex tasks (e.g., GPQA) or smaller models. These results emphasize that naively incorporating a few examples into prompts does not guarantee consistent gains, and that ReFeri is more robust and scalable. Finally, we note that the application of prompt-based approach could be limited due to inherent input context-window length.
Appendix C Qualitative Examples
@@
In this section, we present qualitative examples to further analyze the proposed ReFeri. For better readability, we only present the examples from MATH500, GPQA, and HotpotQA. All the responses are generated by GPT-4o-mini.
C.1 Token level analysis
@@
According to the Hardy-Weinberg equation p^2 + 2pq + q^2 = 1, where p is the frequency of dominant allele frequency, q is the frequency of recessive allele frequency, p^2 is the frequency of the homozygous dominant allele, q^2 is the frequency of the recessive allele, and 2pq is the frequency of the heterozygous allele. Given that q^2=1/400, hence, q=0.05 and p=1-q=0.95. The frequency of the heterozygous allele is 2pq=2*0.05*0.95=38/400.
@@
C.2 Response level analysis
For each response selected by ReFeri, we compute the Eq. (10) with $\beta=1$ for both best and worst response. The value in parentheses is the final selection score. The highest score among all candidates corresponds to the best response, and the lowest score represents the worst response. As shown in below examples, the selected response by ReFeri has more accurate reasoning. For instance, in the example of MATH-500, the best-scoring response anchor their reasoning in a coordinate system and follow clean logic. The worst response, on the other hand, ignore spatial cues, misapply subtraction, and over-complicate with lengthy and internally inconsistent steps. Similarly, in GPQA, the best response reasonably combine the results from the two analyses to arrive at the correct answer. The worst response, on the other hand, appears to rely primarily on the idea that “the heavy branching in (A) is consistent with the splitting observed” without making a clear connection to the analyzed results. Thus, one response integrates the data, while the other reduces it to a vague notion of complexity. In HotpotQA, selected answer are consistent with HotpotQA’s multi-hop requirements by making intermediate hops explicit and factually correct, while rejected answer provide unsupported single-hop claims.
@@ C.2 Response level analysis
[…]
@@
[…]
@@
[…]
@@
Appendix D Societal Impact
ReFeri provides a training-free method for selecting promising outputs from LLMs. This makes it particularly valuable in scenarios where labeled data is scarce or where model fine-tuning is impractical such as personalization tasks with limited access to user data, or applications in emerging domains where predefined labels are unavailable. In addition, ReFeri reduces the barrier to deploying LLMs in real-world settings without additional supervision. This may contribute to broader and more efficient adoption of LLMs in resource-constrained environments.
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ system Please reason step by step, and put your final answer within . ————————————————– user Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac{1}{3}$ of the distance. Each hop tires him out so that he continues to hop $\frac{1}{3}$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction.
Let’s think step by step Kevin hops $1/3$ of the remaining distance with every hop. His first hop takes $1/3$ closer. For his second hop, he has $2/3$ left to travel, so he hops forward $(2/3)(1/3)$. For his third hop, he has $(2/3)^{2}$ left to travel, so he hops forward $(2/3)^{2}(1/3)$. In general, Kevin hops forward $(2/3)^{k-1}(1/3)$ on his $k$th hop. We want to find how far he has hopped after five hops. This is a finite geometric series with first term $1/3$, common ratio $2/3$, and five terms. Thus, Kevin has hopped $\frac{\frac{1}{3}\left(1-\left(\frac{2}{3}\right)^{5}\right)}{1-\frac{2}{3}}=\boxed{\frac{211}{243}}$. The answer is 211243
Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$ ”’ Few-shot CoT prompt on MATH500
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ system Please reason step by step, and put your final answer within . ————————————————– user Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$ ”’ Zero-shot CoT prompt on MATH500
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ I have generated the following responses to the question: Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$
Response 0: response0
Response 4: response4
Evaluate these responses. Select the most consistent response based on majority consensus. Start your answer with "The most consistent response is Response X" (without quotes). ”’ Prompt for USC
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ system Please reason step by step, and put your final answer within . ————————————————– user Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac{1}{3}$ of the distance. Each hop tires him out so that he continues to hop $\frac{1}{3}$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction. ”’ Prompt for LEAP mistakes
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Question: question Generated Reasoning: response
Generated Answer: generated_answer
Correct Reasoning: correct_reasoning
Correct Answer: correct_answer
Instruction: Conduct a thorough analysis of the generated answer in comparison to the correct answer. Also observe how the generated reasoning differs from the correct reasoning. Identify any discrepancies, misunderstandings, or errors. Provide clear insights, principles, or guidelines that can be derived from this analysis to improve future responses. We are not focused on this one data point, but rather on the general principle.
Reasoning: <discuss why the generated answer is wrong> Insights: <what principle should be looked at carefully to improve the performance in the future>
”’ Prompt for LEAP low-level principles
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Low-level principles: low_level_principles
Create a list of *unique* and insightful principles to improve future responses based on the analysis above. Focus on capturing the essence of the feedback while eliminating redundancies. Ensure that each point is clear, concise, and directly derived from the introspection results. Create a numbered list of principles. Leave specific details in place. Limit to at most 8 principles.
List of Principles: ”’ Prompt for LEAP high-level principles
{listing}
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ system Please reason step by step, and put your final answer within . ————————————————– user Please carefully note the following principles:
Principles: 1. **Meticulous Verification**: Always verify each step in algebraic processes to prevent errors that can lead to incorrect conclusions.
8. **Continuous Learning and Adaptation**: Stay open to learning from mistakes and adapting methods to improve future problem-solving approaches.
Let’s think step by step Kevin hops $1/3$ of the remaining distance with every hop. His first hop takes $1/3$ closer. …
Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$ ”’ Prompt for LEAP generations
{listing}
@@ Appendix D Societal Impact
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac{1}{3}$ of the distance. Each hop tires him out so that he continues to hop $\frac{1}{3}$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction.
Let’s think step by step Kevin hops $1/3$ of the remaining distance with every hop. His first hop takes $1/3$ closer. …
I have generated the following responses to the question: Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$
Response 0: response0
Response 4: response4
@@
Evaluate these responses. Select the most consistent response based on majority consensus. Start your answer with "The most consistent response is Response X" (without quotes). ”’ Prompt for USC-w/ Fewshot
{listing}
@@
[!ht] {minted}[fontsize=, frame=single, breaklines]python f”’ system Your job is selecting the most accurate response among multiple candidates. You will receive a question and several candidate answers labeled candidate1, candidate2, etc. Please summarize the debate very briefly and then conclude which single candidate is the most plausible. Output exactly in this format: Summary: <brief summary> Conclusion: candidate<number> Remember to choose only one candidate as the final answer. ————————————————– user Please reason step by step, and put your final answer within .
The below examples are well-constructed gold question and answer pairs for the same task.
Let’s think step by step Kevin hops $1/3$ of the remaining distance with every hop. His first hop takes $1/3$ closer. …
Now, let’s select the most proper answer for the given question Question: Convert the point $(0,3)$ in rectangular coordinates to polar coordinates. Enter your answer in the form $(r,\theta),$ where $r>0$ and $0\leq\theta<2\pi.$ candidate1: response 0 … candidate5: response 4 ”’ Prompt for LLM-as-Judge

### v2

@@ B.1 Additional comparison with few-shot prompting-based selection methods
Notably, we observe that both prompt-based selection methods, USC and LLM-as-Judge, are highly sensitive to the order of candidate responses. In our experiments, USC frequently selections were made from the first two responses regardless of correctness; on multiple choice question tasks this pattern is less extreme but skew toward early positions is still visible. Moreover, since USC requires explicit answer extraction, tasks such as MuSR-ta revealed many failure cases (e.g., over 40% failures in Table 6), further highlighting its fragility. This highlights a critical weakness in prompt-based selection: the output is often determined more by position than content. Based on these observations, we conducted additional experiments where we randomly rearranged the order of candidate responses. Indeed, we observed this issue in Table 7; on GPQA, for example, USC’s accuracy varied notably across different permutations (e.g., 46.0 $\rightarrow$ 41.9), demonstrating its sensitivity to presentation order. In contrast, our approach mitigates such ordering artifacts by decoupling few-shot demonstrations from the selection prompt and using them only for scoring. Furthermore, LLM-as-Judge does not perform reliably on more complex tasks (e.g., GPQA showing a noticeable accuracy degradation compared to other methods). These results emphasize that naively incorporating a few examples into prompts does not guarantee consistent gains, and that ReFeri is more robust and scalable. Finally, we note that the application of prompt-based approach could be limited due to inherent input context-window length.
B.2 Application ReFeri to Zero-shot response
As shown in Table 1, Zero-shot CoT often achieves higher accuracy than Few-shot CoT, reflecting the intrinsic knowledge of the model. However, as described in Eq. 3, ReFeri is also applicable to selecting reasoning paths of Zero-shot CoT, although we primarily apply it to Few-shot CoT since it usually yields better reasoning paths (Table 1). With the experiments in Table 8, we verify that applying ReFeri to Zero-shot CoT yields improvements. These results further suggest that the few-shot exemplars in ReFeri mainly function as a post-hoc validation pipeline, rather than as generation guidance as in conventional Few-shot CoT. Also, this effectiveness of ReFeri under decoupling between generation and selection suggests a robust alternative to conventional few-shot prompting strategies, particularly in settings where few-shot examples are ineffective with LLMs.
Table 8: Performance comparison between Zero-shot and ReFeri under zero-shot setting.
Models Methods MATH500 GPQA MuSR-ta
GPT-4o-mini Zero-shot 76.4 43.0 56.2
ReFeri 78.2 43.9 58.8
GPT-4o Zero-shot 77.5 48.8 66.6
ReFeri 80.8 54.0 69.6
Llama-3.1-8B-It Zero-shot 44.2 21.6 39.6
ReFeri 50.8 24.2 41.2
B.3 Multiple run
Table 9: Multiple runs of GPT-4o-mini across methods (Few-shot, CoT-WP, ReFeri).
Methods MATH500 GPQA MuSR-ta Avg
Run 1 Few-shot 74.1 41.0 77.8 64.3
CoT-WP 76.0 43.9 78.4 66.1
ReFeri 77.2 43.4 82.8 67.8
Run 2 Few-shot 74.6 42.9 77.0 64.8
CoT-WP 76.4 40.4 82.8 66.5
ReFeri 78.4 43.9 82.8 68.4
Run 3 Few-shot 74.9 41.8 75.8 64.2
CoT-WP 76.6 45.0 80.0 67.2
ReFeri 79.2 44.9 80.8 68.3
Variance Few-shot 0.11 0.64 0.68 0.07
CoT-WP 0.06 4.86 5.02 0.21
ReFeri 0.73 0.39 1.07 0.09
To assess the stability of ReFeri, we conduct multiple runs of GPT-4o-mini and reported in Table 9. By design, our method is deterministic for a given set of candidate responses. However, since the candidates themselves are subject to sampling stochasticity, we perform multiple runs to evaluate the consistency of the performance gains. In these runs, we compare ReFeri with the few-shot and CoT-WP baselines. As the table illustrates, ReFeri consistently outperforms both alternatives in all runs. Notably, ReFeri also proves to be significantly more stable. Whereas CoT-WP shows considerable variance in tasks like GPQA (4.9) and MuSR-ta (5.0), ReFeri maintains low variance, indicating that its improvements are reliable despite the stochastic nature of candidate generation. This demonstrates the robustness of ReFeri to sampling variations.
B.4 More results with different few-shot examples
Table 10: Results of GPT-4o-mini across different few-shot examples and ReFeri.
Methods MATH500 GPQA MuSR-ta
Few-shot 1 75.2 41.3 77.0
ReFeri 1 77.8 41.9 82.8
Few-shot 2 74.5 41.5 57.8
ReFeri 2 79.0 43.4 59.2
Few-shot 3 75.0 38.9 60.1
ReFeri 3 77.8 41.9 62.8
In addition to the results reported in Section 3.3, we provide extended experiments in Table 10 including MuSR-ta benchmark. Interestingly, MuSR-ta once again highlights the importance of example quality; when synthesizing new data according to (Sprague et al., 2024) to use as few-shot examples, baseline accuracy significantly degrades. Nevertheless, ReFeri demonstrates consistent performance improvements and confirming the robustness.
Table 11: Performance of ReFeri in weak few-shot settings.
Models Methods MATH500 GPQA MuSR-ta
GPT-4o-mini Few-shot 73.7 41.1 57.5
ReFeri 76.4 43.4 58.0
GPT-4o Few-shot 75.2 46.1 71.0
ReFeri 79.0 47.0 75.6
Llama-3.1-8B-It Few-shot 39.5 26.6 38.7
ReFeri 47.6 33.3 41.6
We believe that constructing accurate few-shot examples is a minimal effort that one should invest to guide LLMs (even humans) toward a proper behavior for a target task. Still, to evaluate robustness, we conducted experiments using intentionally synthesized "weak few-shot" (Table 11) by GPT-4o-mini. Even under these weaker conditions, ReFeri continues to improve performance relative to the Few-shot CoT, confirming that verification remains effective even with suboptimal examples and can provide meaningful gains in more practical, less curated scenarios.
Table 12: Judgment scores (1–10) by GPT-4o for weak fewshot quality.
Judge by GPT-4o (1–10) MATH500 GPQA MuSR-ta
Few-shot 8 8 8
Low quality 3 5 4
To verify the degradation, we asked GPT-4o to evaluate the quality of the original set and the weak example set. The evaluation was conducted in random order, and information about each set was not provided to avoid bias. As shown in Table 12, the weak set consistently received significantly lower scores (3–5 points) compared to the original examples (8 points). This further demonstrates that ReFeri maintains its effectiveness even when the quality of the provided examples is low. See list 14-D.2 for exact prompts used in the generation and evaluation assessment.
B.5 More results on generation/evaluation prompts
Table 13: Ablation on generation/evaluation prompts.
Gen Prompt Eval Prompt MATH500 GPQA MuSR-ta
Few-shot ReFeri Few-shot ReFeri Few-shot ReFeri
Orig Orig 75.2 77.8 41.3 41.9 77.0 82.8
Plan 75.2 78.0 41.3 42.4 77.0 82.8
Role 75.2 77.8 41.3 41.9 77.0 82.4
Plan Plan 74.6 78.2 42.6 47.5 77.0 82.4
Orig 74.6 78.4 42.6 47.5 77.0 82.4
Role Role 74.5 78.2 43.5 47.5 75.8 81.6
Orig 74.5 78.2 43.5 47.0 75.8 81.6
In addition to the prompt style (see 16, 17) ablation study reported in Section 3.3, Table 13 extends the results to include MuSR-ta. As mentioned above, ReFeri demonstrates stable performance across various combinations of generation and evaluation prompts (orig, plan, role), indicating robustness to changes in prompt style. The accuracy of responses generated by Few-shot CoT varies depending on the generation style, but ReFeri consistently shows improved performance across all configurations.
B.6 Full results with different estimation models
Table 14: Full results with different estimation models across three benchmarks.
(a) MATH500
Estimation GPT-4o-mini GPT-4o Llama-3.1-8B-It Avg
Llama-3.2-1B 78.0 77.6 51.4 69.0
Llama-3.1-8B 77.8 77.8 51.2 68.9
Qwen-2.5-7B 78.8 79.2 52.0 70.0
Llama-3.1-70B 77.8 77.6 53.6 69.7
(b) GPQA
Estimation GPT-4o-mini GPT-4o Llama-3.1-8B-It Avg
Llama-3.2-1B 43.9 50.5 33.8 42.7
Llama-3.1-8B 41.9 51.5 35.4 42.9
Qwen-2.5-7B 41.4 50.5 34.3 42.1
Llama-3.1-70B 42.4 53.5 34.8 43.6
(c) MuSR-ta
Estimation GPT-4o-mini GPT-4o Llama-3.1-8B-It Avg
Llama-3.2-1B 83.2 90.8 80.0 84.7
Llama-3.1-8B 82.8 90.4 79.6 84.3
Qwen-2.5-7B 82.0 90.8 81.6 84.8
Llama-3.1-70B 83.6 91.2 81.6 85.5
Table 15: Computational cost. Evaluation cost of GPT-4o-mini. Costs are measured in actual processing time (seconds) per instance on a single GPU using the same model configuration.
Size Methods MATH500 GPQA MuSR-ta
(Acc. / Time) (Acc. / Time) (Acc. / Time)
1B USC 75.0 / 0.6 44.9 / 0.1 75.6 / 0.7
CoT-WP 76.0 / 1.5 43.4 / 2.0 77.6 / 5.0
ReFeri 78.0 / 3.0 43.9 / 4.0 83.2 / 8.0
8B USC 77.8 / 3.7 46.0 / 3.7 74.4 / 3.9
CoT-WP 77.8 / 8.3 42.4 / 11.0 78.8 / 25.6
ReFeri 77.8 / 16.6 41.9 / 22.1 82.8 / 41.8
Table 14 provides full results for all estimation model combinations of MATH500, GPQA and MuSR-ta. This complements the average performance across different generation LLMs (GPT-4o-mini, GPT-4o, and LLaMA3.1-8B) shown in Figure 4. Across all three tasks, ReFeri shows consistent performance gains regardless of the estimation model used, emphasizing its robustness. There are some model-specific trends; for example, smaller models (LLaMA-3.2-1B) perform competitively on (relatively) simple tasks like MATH500, as discussed in Section 3.2.
Moreover, we provide further results by including MuSR-ta in a cost-accuracy analysis (Table 15), which complements the discussion in Sec. 3.3. On this benchmark, name with 1B estimator achieves 83.2% accuracy while requiring only 8s per query, but clearly outperforms the robust 8B CoT-WP baseline, which achieves 78.8% but consumes more than three times the latency (25.6s). This result illustrates that ReFeri with a smaller estimator can still effectively utilizes a few-shot examples to provide robust validation at a much lower cost, making it particularly attractive for scenarios where latency and resource budgets are critical.
B.7 Additional ablation
Table 16: Additional ablation study on GPT-4o-mini
Methods MuSR-ta MuSR-op GPQA MATH500 DROP HotpotQA MMLU-PRO Avg.
(Acc.) (Acc.) (Acc.) (Acc.) (EM / F1) (EM / F1) (Acc.)
No replace (full) 82.8 60.2 42.4 78.0 78.4 / 84.2 36.2 / 48.0 65.0 63.3
No replace 82.4 60.2 42.9 77.6 78.4 / 84.1 35.8 / 47.6 64.7 63.1
ReFeri (Full) 82.8 61.3 42.4 77.8 79.6 / 85.3 35.8 / 47.9 65.0 63.5
ReFeri 82.8 61.3 41.9 77.8 79.2 / 84.9 36.2 / 48.0 64.9 63.4
Here, we conduct the additional experiments to provide comprehensive ablation study for ReFeri. We first evaluate the Full variant (Eq. 8), which generally achieves the strongest results across benchmarks (Table 16). This is expected, as using the complete set of examples provides the most faithful estimate of backward consistency. However, as discussed in Sec. 2.2, the computational overhead increases linearly with the number of few-shot examples, which renders the Full variant less appealing for large-scale or resource-constrained scenarios.
To further analyze this trade-off, we examine the effectiveness of the proposed prompt replacement (Eq. 7) for better estimation of backward score. To this end, we consider a simplified variant of our backward score, termed No replace, where each few-shot example $\mathbf{x}_{i}=(q_{i},a_{i})$ is evaluated in a one-shot manner using the test query $\widetilde{q}$ and the candidate response $r_{k}$ as additional context. Specifically, this variant modifies the backward score in Eq. 8 by replacing the leave-one-out prompt $\widetilde{\mathbf{X}}_{i}$ with a single pair $\mathbf{y}_{k}=(\widetilde{q},r_{k})$:
We note that, as in our main method, a cost-efficient variant can be obtained by incorporating the $i^{\dagger}$ selection strategy (Eq. 9), which adaptively chooses the most relevant exemplar to the test query.
$S^{\prime}_{\texttt{Back}}(r_{k}):=\log P(a_{i^{\dagger}}\mid q_{i^{\dagger}},\widetilde{q},r_{k})-\log P(a_{i^{\dagger}}\mid q_{i^{\dagger}}),$
Appendix C Usage of AI assistants
This paper used AI-based writing aids to improve sentence structure, correct grammar, and improve readability. These tools were applied only to language refinement and did not affect the advancement of technical content, research methodology, or experimental analysis. All scientific ideas, results, and conclusions were conceived and written entirely by researchers. The use of AI aids was limited to editorial purposes and did not impair the originality or intellectual contribution of the work.
Appendix D Qualitative Examples
@@
In this section, we present qualitative examples to further analyze the proposed ReFeri. For better readability, we only present the examples from MATH500, GPQA, and HotpotQA. All the responses are generated by GPT-4o-mini, and we use the ReFeri (Full) variant for illustration to provide the clearest comparisons.
D.1 Token level analysis
@@
According to the Hardy-Weinberg equation pˆ2 + 2pq + qˆ2 = 1, where p is the frequency of dominant allele frequency, q is the frequency of recessive allele frequency, pˆ2 is the frequency of the homozygous dominant allele, qˆ2 is the frequency of the recessive allele, and 2pq is the frequency of the heterozygous allele. Given that qˆ2=1/400, hence, q=0.05 and p=1-q=0.95. The frequency of the heterozygous allele is 2pq=2*0.05*0.95=38/400.
@@
D.2 Response level analysis
For each response selected by ReFeri, we compute the Eq. 11 for both best and worst response. The value in parentheses is the final selection score. The highest score among all candidates corresponds to the best response, and the lowest score represents the worst response. As shown in below examples, the selected response by ReFeri has more accurate reasoning. For instance, in the example of MATH-500, the best-scoring response anchor their reasoning in a coordinate system and follow clean logic. The worst response, on the other hand, ignore spatial cues, misapply subtraction, and over-complicate with lengthy and internally inconsistent steps. Similarly, in GPQA, the best response reasonably combine the results from the two analyses to arrive at the correct answer. The worst response, on the other hand, appears to rely primarily on the idea that “the heavy branching in (A) is consistent with the splitting observed” without making a clear connection to the analyzed results. Thus, one response integrates the data, while the other reduces it to a vague notion of complexity. In HotpotQA, selected answer are consistent with HotpotQA’s multi-hop requirements by making intermediate hops explicit and factually correct, while rejected answer provide unsupported single-hop claims.
@@ D.2 Response level analysis
2. Determine the area of the upper triangle: The shaded area in the upper left can be identified as a right triangle, located at the top left corner of the rectangle. The base of this triangle is the height of the rectangle (4 squares), and the height is the width of the rectangle (2 squares). The area of a triangle is computed using the formula: $\text{Area of triangle}=\frac{1}{2}\times\text{base}\times\text{height}$ For this triangle: $\text{Area}=\frac{1}{2}\times 2\times 4=\frac{1}{2}\times 8=4\text{ square units}$ 3. Determine the area of the other shaded region: The other shaded area is the rectangle that extends across the top down to the base of the triangle. The rectangle also covers one of the 1x1 squares fully, specifically, it will cover the upper two squares fully which are not part of the triangle. This gives us a square area of $3\text{ (width)}\times 4\text{ (height)}-\text{Area of triangle}=12-4$. Thus, the shaded area for consideration is: $12-4=8\text{ square units}$ 4. Calculate the total shaded area: Adding these two areas we’ve established gives: - Area of upper triangle: $4$ - Area of square: 88
@@
5. Final answer: Thus, the area of the shaded region is confirmed as: $\boxed{4}$
@@
Thus, the total area of the shaded region is $\boxed{10}$.
@@
f”’
{system}
[…]
{user}
@@
Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac${1}{3}\$ of the distance. Each hop tires him out so that he continues to hop $\frac${1}{3}\$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction.
[…]
Kevin hops \$1/3\$ of the remaining distance with every hop.
His first hop takes \$1/3\$ closer.
For his second hop, he has \$2/3\$ left to travel, so he hops forward \$(2/3)(1/3)\$.
For his third hop, he has \$(2/3)^2\$ left to travel, so he hops forward \$(2/3)^2(1/3)\$.
[…]
We want to find how far he has hopped after five hops.
This is a finite geometric series with first term \$1/3\$, common ratio \$2/3\$, and five terms.
[…]
The answer is $\frac${211}{243}}
[…]
Figure 5: Few-shot CoT prompt on MATH500
f”’
{system}
[…]
{user}
[…]
Figure 6: Zero-shot CoT prompt on MATH500
f”’
[…]
Select the most consistent response based on majority consensus.
@@
Start your answer with "The most consistent response is Response X" (without quotes).
Figure 7: Prompt for USC
f”’
{system}
[…]
{user}
Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac${1}{3}\$ of the distance. Each hop tires him out so that he continues to hop $\frac${1}{3}\$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction.
Figure 8: Prompt for LEAP mistakes
f”’
Question: {question}
Generated Reasoning: {response}
[…]
Insights: <what principle should be looked at carefully to improve the performance in the future>
[…]
Figure 9: Prompt for LEAP low-level principles
f”’
Low-level principles:
{low_level_principles}
[…]
Focus on capturing the essence of the feedback while eliminating redundancies.
Ensure that each point is clear, concise, and directly derived from the introspection results.
Create a numbered list of principles. Leave specific details in place.
Limit to at most 8 principles.
[…]
Figure 10: Prompt for LEAP high-level principles
f”’
{system}
[…]
{user}
Please carefully note the following principles:
[…]
Kevin hops \$1/3\$ of the remaining distance with every hop.
His first hop takes \$1/3\$ closer.
[…]
Figure 11: Prompt for LEAP generations
f”’
Kevin Kangaroo begins hopping on a number line at 0. He wants to get to 1, but he can hop only $\frac${1}{3}\$ of the distance. Each hop tires him out so that he continues to hop $\frac${1}{3}\$ of the remaining distance. How far has he hopped after five hops? Express your answer as a common fraction.
[…]
Kevin hops \$1/3\$ of the remaining distance with every hop.
His first hop takes \$1/3\$ closer.
[…]
Select the most consistent response based on majority consensus.
Start your answer with "The most consistent response is Response X" (without quotes).
Figure 12: Prompt for USC-w/ Fewshot
f”’
{system}
@@
Your job is selecting the most accurate response among multiple candidates. You will receive a question and several candidate answers labeled candidate1, candidate2, etc. Please summarize the debate very briefly and then conclude which single candidate is the most plausible. Output exactly in this format:
Summary: <brief summary>
Conclusion: candidate<number>
Remember to choose only one candidate as the final answer.
{user}
[…]
Kevin hops \$1/3\$ of the remaining distance with every hop.
His first hop takes \$1/3\$ closer.
[…]
candidate1: {response 0}
candidate5: {response 4}
Figure 13: Prompt for LLM-as-Judge
f”’
[…]
’{"answer":"<rewritten weaker answer>"}’
Figure 14: Prompt for generate weak few-shot
[…]
Figure 15: Prompt for judging weak few-shot
f”’
[…]
Figure 16: Prompt for plan-and-sovle on MATH500
f”’
[…]
Figure 17: Prompt for role-playing on MATH500